• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Rikki

    Log entry time 16:05:28 on February18,2002

    Entry number 83070

    keyword=day shift summary

    08:00 Continue with 3D(c)
    run #2614 Total Q=1.2 C

    9:14 MCC can get more beam current (they adjusted injector A1 slit) so we've asked for the max. they can get without trips. 50uA.

    12:25 Finished up 3D(c)
    run #2622 Total Q ~1.9 C. 8.9 m triggers.

    12:39 Begin 3D(e) with 10 uA.
    run #2623

    14:38 MCC wants to do a little beam tuning. 15 minute estimate.
    15:06 Beam back. Continue with 3D(e) running.

    16:00 run #2634 in progress. 0.09 C as of #2633.